I have one more for you Stan, thanks for the last one. Here it goes.
What amount of each mixture, one 95% alcohol and the other 15% alcohol, must be used to make 10 liters of a mixture which is 45% alcohol? What equation or equations will express this problem?
A. 0.95x + 0.15y = 10 0.45(x+y) = 0.45 x 10

B. 0.95x + 0.15x = 0.45
C. 0.95x + 0.15(10 - x)= 0.45 x 10
D. x + y = 10 0.45x + 0.95y = 0.45 x 10

Answer by stanbon(75887)

What amount of each mixture, one 95% alcohol and the other 15% alcohol, must be used to make 10 liters of a mixture which is 45% alcohol? What equation or equations will express this problem?
95% mixture DATA:
amount = x liters ; active ingredient= 0.95x
-----------------------
15% mixture DATA:
amount = 10-x liters ; active ingredient = 0.15(10-x)
------------------------
45% mixture DATA:
amount =10 liters ; active ingredient = 0.45(10)
-------------------------
EQUATION:
active + active = active
0.95% + 0.15(10-x)= 0.45(10)
Looks like "C" is the answer.
Cheers,
Stan H.